# Cooperative Multi-Agent Planning with Adaptive Skill Synthesis

### v1

@@ 1. Analyze the provided script’s effectiveness
elif obs_data.own_unit_type.lower() in [’zealot’, zergling’, baneling’]:
# If there are enemies
if obs_data.enemies:
enemy_in_range = [enemy for enemy in obs_data.enemies if enemy.distance < 1]
attack_actions = [a for a in valid_actions if a >= 6]
if len(enemy_in_range) > 0:
enemy_center = (sum(e.position[0] for e in enemy_in_range) / len(enemy_in_range),
sum(e.position[1] for e in enemy_in_range) / len(enemy_in_range))
else:
enemy_center = (sum(e.position[0] for e in obs_data.enemies) / len(obs_data.enemies),
sum(e.position[1] for e in obs_data.enemies) / len(obs_data.enemies))
if obs_data.allies:
@@
# Check if any melee ally, if so and last action is not attack, move to center of melee allies
@@
melee_ally = [ally for ally in obs_data.allies if ally.unit_type.lower() in [’zealot’, zergling’, baneling’]]
if melee_ally:
# Move to center of melee allies
ally_x = sum(ally.position[0] for ally in melee_ally) / len(melee_ally) / 2
ally_y = sum(ally.position[1] for ally in melee_ally) / len(melee_ally) / 2
safe_x = ally_x
safe_y = ally_y
distance = (safe_x ** 2 + safe_y ** 2) ** 0.5
criterion = 2/obs_data.own_sight_range
if len(attack_actions) == 0 or distance > 0.5:
target_angle = math.atan2(enemy_center[1], enemy_center[0])
safe_angle = math.atan2(ally_y, ally_x)
angle_diff = abs(target_angle - safe_angle)
if distance > criterion and (math.pi/9 < angle_diff < 17*math.pi/9 or distance > 0.5):
path_action = find_path(obs_data, safe_x, safe_y)
if path_action:
return path_action
# Enhanced cluster detection with dynamic radius
enemy_clusters = {}
cluster_centers = {}
for enemy in obs_data.enemies:
nearby_enemies = []
center_x, center_y = enemy.position[0], enemy.position[1]
# Dynamic cluster radius based on unit type
@@
cluster_radius = 0.3 if obs_data.own_unit_type.lower() == baneling else 0.2
for other in obs_data.enemies:
distance = ((other.position[0] - enemy.position[0])**2 +
(other.position[1] - enemy.position[1])**2)**0.5
if distance <= cluster_radius:
nearby_enemies.append(other)
center_x += other.position[0]
center_y += other.position[1]
if nearby_enemies:
center_x /= len(nearby_enemies)
center_y /= len(nearby_enemies)
enemy_clusters[enemy.id] = len(nearby_enemies)
cluster_centers[enemy.id] = (center_x, center_y)
# Enhanced target scoring with tactical considerations
target_scores = {}
for enemy in obs_data.enemies:
base_score = score_target(enemy)
# Enhanced cluster bonus for splash damage
if obs_data.own_unit_type.lower() == baneling’:
cluster_bonus = 1.5 ** enemy_clusters[enemy.id]
else:
cluster_bonus = 1.2 ** enemy_clusters[enemy.id]
# Calculate final score with all factors
target_scores[enemy.id] = (base_score + cluster_bonus)
# Check if there are same max score targets
max_score = max(target_scores.values())
max_score_targets = [enemy for enemy in obs_data.enemies
if target_scores[enemy.id] >= max_score] # Allow for close scores
if len(max_score_targets) > 1:
# Choose target balancing distance and cluster potential
best_target = min(max_score_targets,
key=lambda x: x.distance)
else:
best_target = max_score_targets[0]
if best_target.can_attack:
return attack(best_target.id)
else:
# Move to the target
dx = best_target.position[0]
dy = best_target.position[1]
@@
path_action = find_path(obs_data, dx, dy, target_type=best_target.unit_type.lower())
if path_action:
return path_action
elif attack_actions:
attackable_enemies = [enemy for enemy in obs_data.enemies if enemy.can_attack]
if obs_data.last_action in attack_actions:
return obs_data.last_action
if attackable_enemies:
return attack(min(attackable_enemies, key=lambda e: e.distance).id)
return random.choice(attack_actions)
# If there are no enemies
else:
# If there are only allies
if obs_data.allies:
# Improved melee group formation
melee_allies = [ally for ally in obs_data.allies
if ally.unit_type.lower() in [’zealot’, zergling’, baneling’]]
if melee_allies:
spacing = 0.1 if obs_data.own_unit_type.lower() == baneling else 0.05
# Dynamic group positioning
center_x = sum(ally.position[0] for ally in melee_allies) / len(melee_allies)
center_y = sum(ally.position[1] for ally in melee_allies) / len(melee_allies)
# Calculate spread from center
max_spread = max(((ally.position[0] - center_x)**2 +
(ally.position[1] - center_y)**2)**0.5
for ally in melee_allies)
own_distance = ((center_x)**2 + (center_y)**2)**0.5
if own_distance > spacing or max_spread > 0.1:
# Move toward center while maintaining minimum spacing
adjusted_x = center_x * 0.85 # Slight offset to prevent overcrowding
adjusted_y = center_y * 0.85
path_action = find_path(obs_data, adjusted_x, adjusted_y)
if path_action:
return path_action
else:
ally_x = sum(ally.position[0] for ally in obs_data.allies) / len(obs_data.allies)
ally_y = sum(ally.position[1] for ally in obs_data.allies) / len(obs_data.allies)
distance = (ally_x ** 2 + ally_y ** 2) ** 0.5
if distance > 0.05:
dx = ally_x
dy = ally_y
path_action = find_path(obs_data, dx, dy)
if path_action:
return path_action
# Ranged units control logic
else:
attack_actions = [a for a in valid_actions if a >= 6]
# If there are enemies
if obs_data.enemies:
# If there are both allies and enemies
# Calculate retreat position
enemy_in_range = [enemy for enemy in obs_data.enemies if enemy.distance < 1]
if len(enemy_in_range) > 0:
enemy_center = (sum(e.position[0] for e in enemy_in_range) / len(enemy_in_range),
sum(e.position[1] for e in enemy_in_range) / len(enemy_in_range))
else:
enemy_center = (sum(e.position[0] for e in obs_data.enemies) / len(obs_data.enemies),
sum(e.position[1] for e in obs_data.enemies) / len(obs_data.enemies))
if obs_data.allies:
@@
# Check if any melee ally, if so and last action is not attack, move to center of melee allies
@@
melee_ally = [ally for ally in obs_data.allies if ally.unit_type.lower() in [’zealot’, zergling’, baneling’]]
@@
melee_enemy = [enemy for enemy in enemy_in_range if enemy.unit_type.lower() in [’zealot’, zergling’, baneling’]]
if melee_ally:
# Move to center of melee allies
ally_x = sum(ally.position[0] for ally in melee_ally) / len(melee_ally)
ally_y = sum(ally.position[1] for ally in melee_ally) / len(melee_ally)
else:
ally_x = sum(ally.position[0] for ally in obs_data.allies) / len(obs_data.allies) / 2
ally_y = sum(ally.position[1] for ally in obs_data.allies) / len(obs_data.allies) / 2
dx = ally_x - enemy_center[0]
dy = ally_y - enemy_center[1]
safe_x = ally_x
safe_y = ally_y
if melee_ally:
safe_x = safe_x + (dx / abs(dx)) * 2/obs_data.own_sight_range if dx != 0 else safe_x
safe_y = safe_y + (dy / abs(dy)) * 2/obs_data.own_sight_range if dy != 0 else safe_y
melee_threaten = False
if melee_enemy:
closest_melee_enemy = min(melee_enemy, key=lambda x: x.distance)
if closest_melee_enemy.distance <= 4/obs_data.own_sight_range:
melee_threaten = True
dx = safe_x - closest_melee_enemy.position[0]
dy = safe_y - closest_melee_enemy.position[1]
safe_x = safe_x + (dx / abs(dx)) * 1/obs_data.own_sight_range if dx != 0 else safe_x
safe_y = safe_y + (dy / abs(dy)) * 1/obs_data.own_sight_range if dy != 0 else safe_y
distance = (safe_x ** 2 + safe_y ** 2) ** 0.5
criterion = 4/obs_data.own_sight_range
if obs_data.last_action >= 6 or len(attack_actions) == 0 or distance > 0.9:
target_angle = math.atan2(enemy_center[1], enemy_center[0])
safe_angle = math.atan2(ally_y, ally_x)
angle_diff = abs(target_angle - safe_angle)
if distance > criterion and ((math.pi/9 < angle_diff < 17*math.pi/9) or melee_threaten or distance > 0.9):
path_action = find_path(obs_data, safe_x, safe_y)
if path_action:
return path_action
# Focus fire logic
# Count how many allies are attacking each enemy
target_counts = {}
for ally in obs_data.allies:
if ally.last_action >= 6:
target_id = ally.last_action - 6
target_counts[target_id] = target_counts.get(target_id, 0) + 1
@@
# Distance to safe point of each enemy affect target choosing
enemy_safe_distance = {enemy.id: ((enemy.position[0] - safe_x) ** 2 + (enemy.position[1] - safe_y) ** 2) ** 0.5 for enemy in obs_data.enemies}
@@
# Find best target combining focus fire and threat scoring
target_scores = {enemy.id: score_target(enemy) for enemy in obs_data.enemies}
for target_id, count in target_counts.items():
if target_id in target_scores:
target_scores[target_id] += count * 0.5
for target_id, scores in target_scores.items():
target_scores[target_id] = scores * (1 - enemy_safe_distance[target_id] * 0.3)
best_target_id = max(target_scores.items(), key=lambda x: x[1])[0]
best_target = next(enemy for enemy in obs_data.enemies if enemy.id == best_target_id)
if best_target.can_attack:
return attack(best_target_id)
else:
# Best target is not in shoot range, move to target
dx = best_target.position[0]
dy = best_target.position[1]
@@
# Only move to target if its direction is not conflicting with the safe point
# Check if target direction aligns with safe point direction
target_angle = math.atan2(dy, dx)
safe_angle = math.atan2(ally_y, ally_x)
angle_diff = abs(target_angle - safe_angle)
# Only move if angle difference is less than 90 degrees
if angle_diff < math.pi/9 or angle_diff > 17*math.pi/9 or not melee_ally:
if best_target.distance > obs_data.own_shoot_range / obs_data.own_sight_range:
@@
path_action = find_path(obs_data, dx, dy, target_type=best_target.unit_type.lower())
if path_action:
return path_action
if attack_actions:
if obs_data.last_action in attack_actions:
return obs_data.last_action
attackable_enemies = [enemy for enemy in obs_data.enemies if enemy.can_attack]
closest_enemy = min(
[enemy for enemy in attackable_enemies],
key=lambda enemy: enemy.distance,
default=None,
if closest_enemy and closest_enemy.can_attack:
return attack(closest_enemy.id)
return random.choice(attack_actions)
else:
if distance > criterion:
path_action = find_path(obs_data, safe_x, safe_y)
if path_action:
return path_action
# If there are only enemies
else:
# Closest enemy as target
closest_enemy = min(
[enemy for enemy in obs_data.enemies],
key=lambda enemy: enemy.distance,
default=None,
# No allies, kitting melee enemies
@@
if closest_enemy.unit_type.lower() in [’zealot’, zergling’, baneling’]:
@@
if closest_enemy.distance <= 4 / obs_data.own_sight_range and obs_data.last_action >= 6:
enemy_x = sum(e.position[0] for e in obs_data.enemies) / len(obs_data.enemies)
enemy_y = sum(e.position[1] for e in obs_data.enemies) / len(obs_data.enemies)
target_x = - enemy_x
target_y = - enemy_y
g_x = target_x * obs_data.own_sight_range + (obs_data.own_position[0] * 32)
g_y = target_y * obs_data.own_sight_range + (obs_data.own_position[1] * 32)
if not (0 <= g_x <= 32 and 0 <= g_y <= 32):
target_x = (0.5 - obs_data.own_position[0]) * 32 / obs_data.own_sight_range
target_y = (0.5 - obs_data.own_position[1]) * 32 / obs_data.own_sight_range
path_action = find_path(obs_data, target_x, target_y)
if path_action:
return path_action
if closest_enemy.can_attack:
return attack(closest_enemy.id)
else:
# No melee enemies, highest priority enemy as target
target_scores = {enemy.id: score_target(enemy) for enemy in obs_data.enemies}
# Check if there are same max score targets
max_score = max(target_scores.values())
max_score_target_ids = [target_id for target_id, score in target_scores.items() if score == max_score]
max_score_targets = [enemy for enemy in obs_data.enemies if enemy.id in max_score_target_ids]
if len(max_score_targets) > 1:
# Chose the closest target
best_target = min(max_score_targets, key=lambda x: x.distance)
else:
best_target = max_score_targets[0]
if best_target.can_attack:
return attack(best_target.id)
else:
# Best target is not in shoot range, move to target
dx = best_target.position[0]
dy = best_target.position[1]
if best_target.distance > obs_data.own_shoot_range / obs_data.own_sight_range:
@@
path_action = find_path(obs_data, dx, dy, target_type=best_target.unit_type.lower())
if path_action:
return path_action
elif attack_actions:
if obs_data.last_action in attack_actions:
return obs_data.last_action
attackable_enemies = [enemy for enemy in obs_data.enemies if enemy.can_attack]
closest_enemy = min(
[enemy for enemy in attackable_enemies],
key=lambda enemy: enemy.distance,
default=None,
if closest_enemy and closest_enemy.can_attack:
return attack(closest_enemy.id)
return random.choice(attack_actions)
# If there are no enemies
else:
# If there are only allies
if obs_data.allies:
# Check if any melee ally

### v2

@@ 24 Args:
278 elif obs_data.own_unit_type.lower() in [’zealot’, ’zergling’, ’baneling’]:
279 # If there are enemies
280 if obs_data.enemies:
281 enemy_in_range = [enemy for enemy in obs_data.enemies if enemy.distance < 1]
282 attack_actions = [a for a in valid_actions if a >= 6]
283
284 if len(enemy_in_range) > 0:
285 enemy_center = (sum(e.position[0] for e in enemy_in_range) / len(enemy_in_range),
286 sum(e.position[1] for e in enemy_in_range) / len(enemy_in_range))
287 else:
288 enemy_center = (sum(e.position[0] for e in obs_data.enemies) / len(obs_data.enemies),
289 sum(e.position[1] for e in obs_data.enemies) / len(obs_data.enemies))
290 if obs_data.allies:
@@
291 # Check if any melee ally, if so and last action is not attack, move to center of melee allies
@@
292 melee_ally = [ally for ally in obs_data.allies if ally.unit_type.lower() in [’zealot’, ’zergling’, ’baneling’]]
293 if melee_ally:
294 # Move to center of melee allies
295 ally_x = sum(ally.position[0] for ally in melee_ally) / len(melee_ally) / 2
296 ally_y = sum(ally.position[1] for ally in melee_ally) / len(melee_ally) / 2
297
298 safe_x = ally_x
299 safe_y = ally_y
300
301 distance = (safe_x ** 2 + safe_y ** 2) ** 0.5
302 criterion = 2/obs_data.own_sight_range
303 if len(attack_actions) == 0 or distance > 0.5:
304 target_angle = math.atan2(enemy_center[1], enemy_center[0])
305 safe_angle = math.atan2(ally_y, ally_x)
306 angle_diff = abs(target_angle - safe_angle)
307 if distance > criterion and (math.pi/9 < angle_diff < 17*math.pi/9 or distance > 0.5):
308 path_action = find_path(obs_data, safe_x, safe_y)
309 if path_action:
310 return path_action
311
312 # Enhanced cluster detection with dynamic radius
313 enemy_clusters = {}
314 cluster_centers = {}
315 for enemy in obs_data.enemies:
316 nearby_enemies = []
317 center_x, center_y = enemy.position[0], enemy.position[1]
318
319 # Dynamic cluster radius based on unit type
@@
320 cluster_radius = 0.3 if obs_data.own_unit_type.lower() == ’baneling’ else 0.2
321
322 for other in obs_data.enemies:
323 distance = ((other.position[0] - enemy.position[0])**2 +
324 (other.position[1] - enemy.position[1])**2)**0.5
325 if distance <= cluster_radius:
326 nearby_enemies.append(other)
327 center_x += other.position[0]
328 center_y += other.position[1]
329
330 if nearby_enemies:
331 center_x /= len(nearby_enemies)
332 center_y /= len(nearby_enemies)
333
334 enemy_clusters[enemy.id] = len(nearby_enemies)
335 cluster_centers[enemy.id] = (center_x, center_y)
336
337 # Enhanced target scoring with tactical considerations
338 target_scores = {}
339 for enemy in obs_data.enemies:
340 base_score = score_target(enemy)
341
342 # Enhanced cluster bonus for splash damage
343 if obs_data.own_unit_type.lower() == ’baneling’:
344 cluster_bonus = 1.5 ** enemy_clusters[enemy.id]
345 else:
346 cluster_bonus = 1.2 ** enemy_clusters[enemy.id]
347 # Calculate final score with all factors
348 target_scores[enemy.id] = (base_score + cluster_bonus)
349
350 # Check if there are same max score targets
351 max_score = max(target_scores.values())
352 max_score_targets = [enemy for enemy in obs_data.enemies
353 if target_scores[enemy.id] >= max_score] # Allow for close scores
354 if len(max_score_targets) > 1:
355 # Choose target balancing distance and cluster potential
356 best_target = min(max_score_targets,
357 key=lambda x: x.distance)
358 else:
359 best_target = max_score_targets[0]
360
361 if best_target.can_attack:
362 return attack(best_target.id)
363 else:
364 # Move to the target
365 dx = best_target.position[0]
366 dy = best_target.position[1]
@@
367 path_action = find_path(obs_data, dx, dy, target_type=best_target.unit_type.lower())
368 if path_action:
369 return path_action
370 elif attack_actions:
371 attackable_enemies = [enemy for enemy in obs_data.enemies if enemy.can_attack]
372 if obs_data.last_action in attack_actions:
373 return obs_data.last_action
374 if attackable_enemies:
375 return attack(min(attackable_enemies, key=lambda e: e.distance).id)
376 return random.choice(attack_actions)
377
378 # If there are no enemies
379 else:
380 # If there are only allies
381 if obs_data.allies:
382 # Improved melee group formation
383 melee_allies = [ally for ally in obs_data.allies
384 if ally.unit_type.lower() in [’zealot’, ’zergling’, ’baneling’]]
385 if melee_allies:
386 spacing = 0.1 if obs_data.own_unit_type.lower() == ’baneling’ else 0.05
387 # Dynamic group positioning
388 center_x = sum(ally.position[0] for ally in melee_allies) / len(melee_allies)
389 center_y = sum(ally.position[1] for ally in melee_allies) / len(melee_allies)
390
391 # Calculate spread from center
392 max_spread = max(((ally.position[0] - center_x)**2 +
393 (ally.position[1] - center_y)**2)**0.5
394 for ally in melee_allies)
395
396 own_distance = ((center_x)**2 + (center_y)**2)**0.5
397
398 if own_distance > spacing or max_spread > 0.1:
399 # Move toward center while maintaining minimum spacing
400 adjusted_x = center_x * 0.85 # Slight offset to prevent overcrowding
401 adjusted_y = center_y * 0.85
402 path_action = find_path(obs_data, adjusted_x, adjusted_y)
403 if path_action:
404 return path_action
405 else:
406 ally_x = sum(ally.position[0] for ally in obs_data.allies) / len(obs_data.allies)
407 ally_y = sum(ally.position[1] for ally in obs_data.allies) / len(obs_data.allies)
408 distance = (ally_x ** 2 + ally_y ** 2) ** 0.5
409 if distance > 0.05:
410 dx = ally_x
411 dy = ally_y
412 path_action = find_path(obs_data, dx, dy)
413 if path_action:
414 return path_action
415 # Ranged units control logic
416 else:
417 attack_actions = [a for a in valid_actions if a >= 6]
418 # If there are enemies
419 if obs_data.enemies:
420 # If there are both allies and enemies
421 # Calculate retreat position
422 enemy_in_range = [enemy for enemy in obs_data.enemies if enemy.distance < 1]
423 if len(enemy_in_range) > 0:
424 enemy_center = (sum(e.position[0] for e in enemy_in_range) / len(enemy_in_range),
425 sum(e.position[1] for e in enemy_in_range) / len(enemy_in_range))
426 else:
427 enemy_center = (sum(e.position[0] for e in obs_data.enemies) / len(obs_data.enemies),
428 sum(e.position[1] for e in obs_data.enemies) / len(obs_data.enemies))
429 if obs_data.allies:
@@
430 # Check if any melee ally, if so and last action is not attack, move to center of melee allies
@@
431 melee_ally = [ally for ally in obs_data.allies if ally.unit_type.lower() in [’zealot’, ’zergling’, ’baneling’]]
@@
432 melee_enemy = [enemy for enemy in enemy_in_range if enemy.unit_type.lower() in [’zealot’, ’zergling’, ’baneling’]]
433 if melee_ally:
434 # Move to center of melee allies
435 ally_x = sum(ally.position[0] for ally in melee_ally) / len(melee_ally)
436 ally_y = sum(ally.position[1] for ally in melee_ally) / len(melee_ally)
437
438 else:
439 ally_x = sum(ally.position[0] for ally in obs_data.allies) / len(obs_data.allies) / 2
440 ally_y = sum(ally.position[1] for ally in obs_data.allies) / len(obs_data.allies) / 2
441
442 dx = ally_x - enemy_center[0]
443 dy = ally_y - enemy_center[1]
444 safe_x = ally_x
445 safe_y = ally_y
446 if melee_ally:
447 safe_x = safe_x + (dx / abs(dx)) * 2/obs_data.own_sight_range if dx != 0 else safe_x
448 safe_y = safe_y + (dy / abs(dy)) * 2/obs_data.own_sight_range if dy != 0 else safe_y
449 melee_threaten = False
450 if melee_enemy:
451 closest_melee_enemy = min(melee_enemy, key=lambda x: x.distance)
452 if closest_melee_enemy.distance <= 4/obs_data.own_sight_range:
453 melee_threaten = True
454 dx = safe_x - closest_melee_enemy.position[0]
455 dy = safe_y - closest_melee_enemy.position[1]
456 safe_x = safe_x + (dx / abs(dx)) * 1/obs_data.own_sight_range if dx != 0 else safe_x
457 safe_y = safe_y + (dy / abs(dy)) * 1/obs_data.own_sight_range if dy != 0 else safe_y
458
459 distance = (safe_x ** 2 + safe_y ** 2) ** 0.5
460 criterion = 4/obs_data.own_sight_range
461 if obs_data.last_action >= 6 or len(attack_actions) == 0 or distance > 0.9:
462 target_angle = math.atan2(enemy_center[1], enemy_center[0])
463 safe_angle = math.atan2(ally_y, ally_x)
464 angle_diff = abs(target_angle - safe_angle)
465 if distance > criterion and ((math.pi/9 < angle_diff < 17*math.pi/9) or melee_threaten or distance > 0.9):
466 path_action = find_path(obs_data, safe_x, safe_y)
467 if path_action:
468 return path_action
469 # Focus fire logic
470 # Count how many allies are attacking each enemy
471 target_counts = {}
472 for ally in obs_data.allies:
473 if ally.last_action >= 6:
474 target_id = ally.last_action - 6
475 target_counts[target_id] = target_counts.get(target_id, 0) + 1
@@
476 # Distance to safe point of each enemy affect target choosing
477 enemy_safe_distance = {enemy.id: ((enemy.position[0] - safe_x) ** 2 + (enemy.position[1] - safe_y) ** 2) ** 0.5 for enemy in obs_data.enemies}
@@
478 # Find best target combining focus fire and threat scoring
479 target_scores = {enemy.id: score_target(enemy) for enemy in obs_data.enemies}
480 for target_id, count in target_counts.items():
481 if target_id in target_scores:
482 target_scores[target_id] += count * 0.5
483 for target_id, scores in target_scores.items():
484 target_scores[target_id] = scores * (1 - enemy_safe_distance[target_id] * 0.3)
485 best_target_id = max(target_scores.items(), key=lambda x: x[1])[0]
486 best_target = next(enemy for enemy in obs_data.enemies if enemy.id == best_target_id)
487 if best_target.can_attack:
488 return attack(best_target_id)
489 else:
490 # Best target is not in shoot range, move to target
491 dx = best_target.position[0]
492 dy = best_target.position[1]
493
@@
494 # Only move to target if its direction is not conflicting with the safe point
495 # Check if target direction aligns with safe point direction
496 target_angle = math.atan2(dy, dx)
497 safe_angle = math.atan2(ally_y, ally_x)
498 angle_diff = abs(target_angle - safe_angle)
499 # Only move if angle difference is less than 90 degrees
500 if angle_diff < math.pi/9 or angle_diff > 17*math.pi/9 or not melee_ally:
501 if best_target.distance > obs_data.own_shoot_range / obs_data.own_sight_range:
@@
502 path_action = find_path(obs_data, dx, dy, target_type=best_target.unit_type.lower())
503 if path_action:
504 return path_action
505 if attack_actions:
506 if obs_data.last_action in attack_actions:
507 return obs_data.last_action
508 attackable_enemies = [enemy for enemy in obs_data.enemies if enemy.can_attack]
509 closest_enemy = min(
510 [enemy for enemy in attackable_enemies],
511 key=lambda enemy: enemy.distance,
512 default=None,
513 )
514 if closest_enemy and closest_enemy.can_attack:
515 return attack(closest_enemy.id)
516 return random.choice(attack_actions)
517 else:
518 if distance > criterion:
519 path_action = find_path(obs_data, safe_x, safe_y)
520 if path_action:
521 return path_action
522
523 # If there are only enemies
524 else:
525 # Closest enemy as target
526 closest_enemy = min(
527 [enemy for enemy in obs_data.enemies],
528 key=lambda enemy: enemy.distance,
529 default=None,
530 )
531 # No allies, kitting melee enemies
@@
532 if closest_enemy.unit_type.lower() in [’zealot’, ’zergling’, ’baneling’]:
@@
533 if closest_enemy.distance <= 4 / obs_data.own_sight_range and obs_data.last_action >= 6:
534 enemy_x = sum(e.position[0] for e in obs_data.enemies) / len(obs_data.enemies)
535 enemy_y = sum(e.position[1] for e in obs_data.enemies) / len(obs_data.enemies)
536 target_x = - enemy_x
537 target_y = - enemy_y
538
539 g_x = target_x * obs_data.own_sight_range + (obs_data.own_position[0] * 32)
540 g_y = target_y * obs_data.own_sight_range + (obs_data.own_position[1] * 32)
541 if not (0 <= g_x <= 32 and 0 <= g_y <= 32):
542 target_x = (0.5 - obs_data.own_position[0]) * 32 / obs_data.own_sight_range
543 target_y = (0.5 - obs_data.own_position[1]) * 32 / obs_data.own_sight_range
544
545 path_action = find_path(obs_data, target_x, target_y)
546 if path_action:
547 return path_action
548 if closest_enemy.can_attack:
549 return attack(closest_enemy.id)
550 else:
551 # No melee enemies, highest priority enemy as target
552 target_scores = {enemy.id: score_target(enemy) for enemy in obs_data.enemies}
553 # Check if there are same max score targets
554 max_score = max(target_scores.values())
555 max_score_target_ids = [target_id for target_id, score in target_scores.items() if score == max_score]
556 max_score_targets = [enemy for enemy in obs_data.enemies if enemy.id in max_score_target_ids]
557 if len(max_score_targets) > 1:
558 # Chose the closest target
559 best_target = min(max_score_targets, key=lambda x: x.distance)
560 else:
561 best_target = max_score_targets[0]
562 if best_target.can_attack:
563 return attack(best_target.id)
564 else:
565 # Best target is not in shoot range, move to target
566 dx = best_target.position[0]
567 dy = best_target.position[1]
568 if best_target.distance > obs_data.own_shoot_range / obs_data.own_sight_range:
@@
569 path_action = find_path(obs_data, dx, dy, target_type=best_target.unit_type.lower())
570 if path_action:
571 return path_action
572 elif attack_actions:
573 if obs_data.last_action in attack_actions:
574 return obs_data.last_action
575 attackable_enemies = [enemy for enemy in obs_data.enemies if enemy.can_attack]
576 closest_enemy = min(
577 [enemy for enemy in attackable_enemies],
578 key=lambda enemy: enemy.distance,
579 default=None,
580 )
581 if closest_enemy and closest_enemy.can_attack:
582 return attack(closest_enemy.id)
583 return random.choice(attack_actions)
584 # If there are no enemies
585 else:
586 # If there are only allies
587 if obs_data.allies:
588 # Check if any melee ally